# Challenges and Responses in the Practice of Large Language Models

### v1

@@ 2 Software Architecture
1. High computing resource consumption: Large model training requires a lot of computing resources, including high-performance GPUs and a large amount of storage space [3]. This may lead to high training costs and high requirements for hardware resources.
2. Hyperparameter search: The effect of large model training is directly related to the hyperparameter configuration. It is important to search for the best hyperparameters for specific data sets and application scenarios.
3. Data management: Faced with problems such as data diversity, data coverage, data noise, and data quality, large models are prone to underfitting and overfitting problems, resulting in poor performance of the model on new data and hallucinations.
4. Interpretability: The complexity and number of parameters of large models often make their decision-making process opaque, which may lead to difficulties in attribution and tracing. At the same time, this also limits the application of models in scenarios that require high interpretability.
5. Risk control: The training and use of large models may trigger a series of AI safety issues, such as bias, violation, and unfairness. If the training data contains biased, misleading, and toxic information, the model may internalize these biases and lead to corresponding results.
6. Evaluation of large model performance: Use public benchmarks for evaluation. Automated evaluation indicators can be used for some tasks, and manual evaluation can be used for some tasks.
@@ 4 Application Scenarios
However, with the advancement of technology, more and more studies have explored the use of multimodal large models to directly understand and parse tables in documents. This method has shown great potential in complex scenarios and can more intelligently capture the semantics and structural information of tables. Nevertheless, optimizing the document processing process from the source, that is, presenting and submitting the table data in a structured Json format separately during the document preparation stage, is undoubtedly the best practice to improve data processing efficiency and accuracy. This approach not only simplifies the subsequent data extraction and conversion work, but also ensures the consistency and reusability of the data, laying a solid foundation for data analysis and mining.
@@
The knowledge that the RAG model relies on is often discrete and fragmented. In contrast, knowledge graphs organized through graph structures exhibit highly systematic and structured characteristics. Knowledge graph uses graph structure as the storage basis. This design not only promotes the effective integration of knowledge, but also greatly facilitates the knowledge discovery process, which is embodied in a series of advanced functions such as graph storage, graph query, graph search and graph calculation. On support. Once constructed, the knowledge graph forms a self-consistent and comprehensive knowledge system that transcends the boundaries of traditional data storage forms [6], whether it is structured data in databases, semi-structured information in JSON files, or widely distributed Knowledge fragments in unstructured text, as well as multi-modal data, are seamlessly integrated and transformed into interactive graph structures. The advantage of this integration is that users no longer need to care about where the knowledge is stored. They only need to query the graph through natural language or graph query language, and they can freely explore and dig into the required information, which greatly improves knowledge acquisition. efficiency and convenience. Therefore, the knowledge graph is not only an innovation in the way of knowledge representation and storage, but also an important cornerstone for promoting intelligent applications to a higher level.
@@
After the large model identifies the entity, solving the problem of whether USA and America are the same entity mainly involves entity disambiguation and entity linking technologies. The large model only solves one link in the entire requirement. Entity disambiguation refers to solving the problem of polysemy of a word with the same name, such as Apple. Link the identified entity with the entity in the knowledge base (such as Wikipedia, DBpedia, etc.). Determine whether they represent the same object by calculating the similarity between the entity and the entity in the knowledge base (such as vector-based similarity calculation). For example, the entity page and redirect page information in Wikipedia can be used to confirm whether USA and America are linked to the same page. After confirming that USA and America are the same entity, entity normalization is required, that is, all different names pointing to the entity are unified into a standard form (such as ”United States of America”). This helps with subsequent entity association, retrieval, and data analysis.
Question: In the realm of software security, how can knowledge graph technology be leveraged to achieve entity alignment across vulnerability databases? What are the advantages and disadvantages of this approach when compared to large model matching methods?
@@
Advantages: 1. Structured representation: Knowledge graphs represent knowledge in a structured form, making the relationship between entities clearer, more intuitive, and easier to understand and query. 2. Strong interpretability: Compared with the black box characteristics of large models, the alignment process of knowledge graphs is more transparent, and the alignment results can be explained by analyzing entities and relationships. 3. Domain adaptability: In the field of software security, knowledge graphs can make full use of domain expertise and rules to improve the accuracy and pertinence of alignment. 4. Low data dependence: The alignment process of knowledge graphs mainly depends on the structure and relationship of the data itself, and has a low degree of dependence on external training data.
@@
Disadvantages: 1. High construction cost: Building a knowledge graph requires a lot of manpower and time to define entities, relationships, and rules, and to preprocess and clean the data. 2. Poor flexibility: The structure of the knowledge graph is relatively fixed, and it is difficult to quickly adapt to changes and updates in data. In contrast, large models can adapt to new data and tasks through retraining. 3. Dependence on domain knowledge: The construction and alignment process of the knowledge graph requires the participation and guidance of domain experts to ensure accuracy and reliability. This limits its popularity and application scope to a certain extent.
@@
Question: In the field of robotics, does the integration of robots with large models possess significant practical application value?
@@
In the field of robots, especially in the subdivision of domestic robots, the combination of robots and large models has shown extremely broad and practical application value. This combination not only strengthens the robot’s perception and cognitive capabilities, but also greatly improves its multi-modal perception capabilities, making it capable of complex and changeable multi-task scenarios.
@@
Take housekeeping robots as an example, such as advanced systems such as Aloha. They need to handle a series of trivial and detailed tasks from sweeping the floor and folding quilts to cooking and watering flowers. Each task requires the robot to possess different professional knowledge and skills, which places extremely high demands on the robot’s intelligence level. By introducing large models into the design of domestic robots, we can achieve the following significant advantages:
@@
1. Enhanced perception capabilities: The large model can process and analyze data from a variety of sensors, including vision, hearing, touch, etc., thereby giving the robot more comprehensive and accurate perception capabilities. This improvement in multi-modal perception enables robots to better understand and adapt to complex changes in the home environment.
@@
2. Optimized cognitive capabilities: Large models have powerful learning and reasoning [8] capabilities, and can perform knowledge learning and pattern recognition based on massive data. This enables the housekeeping robot to quickly call upon relevant knowledge and formulate and execute reasonable action plans when faced with different tasks. At the same time, large models can also help robots coordinate and optimize tasks to ensure overall work efficiency and effectiveness.
@@
3. Flexible task processing capabilities: With the support of large models, housekeeping robots can handle various tasks more flexibly. Whether it is simple sweeping the floor, folding quilts, or complex tasks such as cooking and watering flowers, the robot can make intelligent judgments and decisions based on the current environment and user needs. In addition, large models can help robots continuously learn and optimize their skills to adapt to changing household needs.
@@
4. Improved user experience: The combination of large models and housekeeping robots not only improves the robot’s work efficiency and accuracy, but also greatly improves the user experience. Users can interact with the robot through natural language, giving instructions or making demands. The robot can accurately understand the user’s intention and give corresponding feedback and execution results. This intelligent interaction method makes housekeeping robots an indispensable assistant and partner in family life.
@@
Question: What are the key technological differences in the stacks employed by various types of AI search, including Perplexity AI, Large-scale model-powered search, AI-powered search solutions from traditional search companies, and AI search startups?
@@
1. Perplexity AI is based on LLMs such as GPT-3.5 and GPT-4, which have powerful language understanding and generation capabilities. Technology integration: Perplexity combines generative AI with search technology, and combines it with Microsoft’s Bing search engine to create a new AI-empowered conversational search engine. Application scenario: Provide direct answers and summaries, and cite relevant sources to support multi-turn conversations and contextual memory capabilities. Optimization strategy: Optimize product performance through continuous iteration, improve user experience, and launch a variety of tool products such as Copilot, Bird SQL, etc.
2. Large-scale model-powered search’s primary focus is on serving as the backbone of a multi-turn dialogue robot. Additionally, it supports the uploading of documents for RAG. Technology integration: Deeply integrate large-scale models with vector databases to alleviate the hallucination problem. Application scenario: Chatbot and Knowledge base Q&A. Optimization strategy: They emphasize the importance of continuously refining and optimizing their foundational models. This includes enhancing the model’s ability to process long contexts, accelerating inference speeds, and implementing other improvements that directly impact model performance and accuracy.
@@
4. AI search startups frequently opt for either open-source or commercialized large-scale models as their foundational platform. They then integrate with search engine APIs, meticulously fine-tuning and optimizing these models to cater to the unique requirements of local markets and individual user preferences. Technology integration: AI search startups integrate advanced technologies such as semantic analysis and knowledge graphs into their foundational models. These integrations enable a deeper understanding of user queries, allowing for more relevant and accurate search results. Application scenarios: You can focus on specific domain or scenarios, such as e-commerce search, academic search, legal search, etc., or you can search in general domain. Optimization strategy: Continuously improve product performance and user experience through technological innovation and iterative optimization, while exploring cooperation opportunities with upstream and downstream enterprises.
@@ 5 Brain Science
On the other hand, the fruitful results of brain science research are profoundly affecting the development trajectory of the field of artificial intelligence. By integrating the profound insights of brain science into the research and development of AI technology, it not only gives artificial intelligence systems capabilities that are closer to human thinking, but also greatly promotes the expansion of the boundaries and performance leaps of AI technology. This interdisciplinary integration not only provides a solid theoretical foundation and source of inspiration for the technological innovation of the AI industry, but also paves the way for the infinite possibilities of future intelligent technology.
@@
2. Memory mechanism: The human brain has a complex and sophisticated memory system, including short-term memory and long-term memory, as well as an efficient memory storage and retrieval mechanism. This biological characteristic provides valuable inspiration for the model architecture in memory processing. By drawing on the brain’s memory mechanism.
3. Multi-brain region collaborative information processing paradigm: Human cognitive functions do not exist in isolation, but rely on close collaboration and information exchange between multiple brain regions to form brain region loops and brain networks for cognitive functions. This multi-brain region collaborative mechanism helps the model’s design ideas when building complex information processing systems. By simulating the functional division of labor and synergy of different brain regions, more complex cognitive tasks can be achieved.
4. Dynamic system perspective of brain-like mechanism: As a highly dynamic system, the internal mechanism of the brain is far beyond the scope of simple electrical signal transmission. Complex phenomena such as the formation and forgetting of memories, the fluctuation and regulation of emotions, etc. often involve complex reactions and regulation of chemical substances. This perspective prompts us to not only focus on the optimization of the computational level when designing the Transformer model, but also explore how to introduce more diversified mechanisms (such as dynamic weight adjustment, emotional computing, etc.) to build a brain-like model that is closer to human intelligence.
5. Energy consumption: The energy consumption of the brain is much lower than that of the Transformer large model, mainly because: Efficient biological components: The energy efficiency of neurons and synapses far exceeds that of electronic components. Parallel and distributed processing: Information processing is highly parallel and distributed, which improves efficiency and reduces energy consumption. Sparse connections: The connections between neurons are sparse, which reduces unnecessary information processing. Adaptive plasticity: It can optimize neural networks based on learning and experience to reduce energy consumption. Evolutionary optimization: Long-term evolution has enabled the brain to develop efficient energy consumption mechanisms. Efficient energy utilization: Relying on glucose for energy, without energy reserves, emphasizing the importance of efficient energy consumption.
@@
The memory of agents can indeed be inspired by brain science. The brain’s memory mechanisms include short-term memory, long-term memory, working memory, etc. These mechanisms provide important references for the memory design of agents. For example, inspired by the brain’s working memory mechanism, the DNC (Differential Neural Computer) in artificial neural networks divides sequence control and memory storage into two modules, improving the ability to handle complex tasks. In addition, the continuous learning mechanism in the brain also provides inspiration for agents to learn another new skill based on mastering one skill.

### v2

@@ 2 Software Architecture
1. High computing resource consumption: big model training requires a lot of computing resources, including high-performance GPUs and a large amount of storage space [3]. This may lead to high training costs and high requirements for hardware resources.
2. Hyperparameter search: The effect of big model training is directly related to the hyperparameter configuration. It is important to search for the best hyperparameters for specific data sets and application scenarios.
3. Data management: Faced with problems such as data diversity, data coverage, data noise, and data quality, big models are prone to underfitting and overfitting problems, resulting in poor performance of the model on new data and hallucinations.
4. Interpretability: The complexity and number of parameters of big models often make their decision-making process opaque, which may lead to difficulties in attribution and tracing. At the same time, this also limits the application of models in scenarios that require high interpretability.
5. Risk control: The training and use of big models may trigger a series of AI safety issues, such as bias, violation, and unfairness. If the training data contains biased, misleading, and toxic information, the model may internalize these biases and lead to corresponding results.
6. Evaluation of big model performance: Use public benchmarks for evaluation. Automated evaluation indicators can be used for some tasks, and manual evaluation can be used for some tasks.
@@ 4 Application Scenarios
However, with the advancement of technology, more and more studies have explored the use of multimodal big models to directly understand and parse tables in documents. This method has shown great potential in complex scenarios and can more intelligently capture the semantics and structural information of tables. Nevertheless, optimizing the document processing process from the source, that is, presenting and submitting the table data in a structured Json format separately during the document preparation stage, is undoubtedly the best practice to improve data processing efficiency and accuracy. This approach not only simplifies the subsequent data extraction and conversion work, but also ensures the consistency and reusability of the data, laying a solid foundation for data analysis and mining.
@@
The knowledge that the RAG model relies on is often discrete and fragmented. In contrast, knowledge graphs organized through graph structures exhibit highly systematic and structured characteristics. Knowledge graph uses graph structure as the storage basis. This design not only promotes the effective integration of knowledge, but also greatly facilitates the knowledge discovery process, which is embodied in a series of advanced functions such as graph storage, graph query, graph search and graph computing. Once constructed, the knowledge graph forms a self-consistent and comprehensive knowledge system that transcends the boundaries of traditional data storage forms [6], whether it is structured data in databases, semi-structured information in JSON files, or widely distributed Knowledge fragments in unstructured text, as well as multi-modal data, are seamlessly integrated and transformed into interactive graph structures. The advantage of this integration is that users no longer need to care about where the knowledge is stored. They only need to query the graph through natural language or graph query language, and they can freely explore and dig into the required information, which greatly improves knowledge acquisition. efficiency and convenience. Therefore, the knowledge graph is not only an innovation in the way of knowledge representation and storage, but also an important cornerstone for promoting intelligent applications to a higher level.
@@
After the big model identifies the entity, solving the problem of whether USA and America are the same entity mainly involves entity disambiguation and entity linking technologies. The big model only solves one link in the entire requirement. Entity disambiguation refers to solving the problem of polysemy of a word with the same name, such as Apple. Link the identified entity with the entity in the knowledge base (such as Wikipedia, DBpedia, etc.). Determine whether they represent the same entity by calculating the similarity between the entity and the entity in the knowledge base (such as vector-based similarity calculation). For example, the entity page and redirect page information in Wikipedia can be used to confirm whether USA and America are linked to the same page. After confirming that USA and America are the same entity, entity normalization is required, that is, all different names pointing to the entity are unified into a standard form (such as ”United States of America”). This helps with subsequent entity association, retrieval, and data analysis.
Question: In the realm of software security, how can knowledge graph technology be leveraged to achieve entity alignment across vulnerability databases? What are the advantages and disadvantages of this approach when compared to big model matching methods?
@@
Advantages: 1. Structured representation: Knowledge graphs represent knowledge in a structured form, making the relationship between entities clearer, more intuitive, and easier to understand and query. 2. Strong interpretability: Compared with the black box characteristics of big models, the alignment process of knowledge graphs is more transparent, and the alignment results can be explained by analyzing entities and relationships. 3. Domain adaptability: In the field of software security, knowledge graphs can make full use of domain expertise and rules to improve the accuracy and pertinence of alignment. 4. Low data dependence: The alignment process of knowledge graphs mainly depends on the structure and relationship of the data itself, and has a low degree of dependence on external training data.
@@
Disadvantages: 1. High construction cost: Building a knowledge graph requires a lot of manpower and time to define entities, relationships, and rules, and to preprocess and clean the data. 2. Poor flexibility: The structure of the knowledge graph is relatively fixed, and it is difficult to quickly adapt to changes and updates in data. In contrast, big models can adapt to new data and tasks through retraining. 3. Dependence on domain knowledge: The construction and alignment process of the knowledge graph requires the participation and guidance of domain experts to ensure accuracy and reliability. This limits its popularity and application scope to a certain extent.
@@
Question: In the field of robotics, does the integration of robots with big models possess significant practical application value?
@@
In the field of robots, the combination of robots and big models has shown extremely broad and practical application value. This combination not only strengthens the robot’s perception and cognitive capabilities, but also greatly improves its multi-modal perception capabilities, making it capable of complex and changeable multi-task scenarios.
@@
Take housekeeping robots as an example, such as advanced systems such as Aloha. They need to handle a series of trivial and detailed tasks from sweeping the floor and folding quilts to cooking and watering flowers. Each task requires the robot to possess different professional knowledge and skills, which places extremely high demands on the robot’s intelligence level. By introducing big models into the design of domestic robots, we can achieve the following significant advantages:
@@
1. Enhanced perception capabilities: The big model can process and analyze data from a variety of sensors, including vision, hearing, touch, etc., thereby giving the robot more comprehensive and accurate perception capabilities. This improvement in multi-modal perception enables robots to better understand and adapt to complex changes in the home environment.
@@
2. Optimized cognitive capabilities: big models have powerful learning and reasoning [8] capabilities, and can perform knowledge learning and pattern recognition based on massive data. This enables the housekeeping robot to quickly call upon relevant knowledge and formulate and execute reasonable action plans when faced with different tasks. At the same time, big models can also help robots coordinate and optimize tasks to ensure overall work efficiency and effectiveness.
@@
3. Flexible task processing capabilities: With the support of big models, housekeeping robots can handle various tasks more flexibly. Whether it is simple sweeping the floor, folding quilts, or complex tasks such as cooking and watering flowers, the robot can make intelligent judgments and decisions based on the current environment and user needs. In addition, big models can help robots continuously learn and optimize their skills to adapt to changing household needs.
@@
4. Improved user experience: The combination of big models and housekeeping robots not only improves the robot’s work efficiency and accuracy, but also greatly improves the user experience. Users can interact with the robot through natural language, giving instructions or making demands. The robot can accurately understand the user’s intention and give corresponding feedback and execution results. This intelligent interaction method makes housekeeping robots an indispensable assistant and partner in family life.
@@
Question: What are the key technological differences in the stacks employed by various types of AI search, including Perplexity AI, Big model-powered search, AI-powered search solutions from traditional search companies, and AI search startups?
@@
1. Perplexity AI is built upon cutting-edge LLMs, including GPT-3.5 and GPT-4, as well as the Bing search engine API, which have powerful language understanding and generation capabilities. Technology integration: Perplexity combines generative AI with search technology, and combines it with Microsoft’s Bing search engine to create a new AI-empowered conversational search engine. Application scenario: Provide direct answers and summaries, and cite relevant sources to support multi-turn conversations and contextual memory capabilities. Optimization strategy: Optimize product performance through continuous iteration, improve user experience, and launch a variety of tool products such as Copilot, Bird SQL, etc.
2. Big model-powered search’s primary focus is on serving as the backbone of a multi-turn dialogue robot. Additionally, it supports the uploading of documents for RAG. Technology integration: Deeply integrate big models with vector databases to alleviate the hallucination problem. Application scenario: Chatbot and Knowledge base Q&A. Optimization strategy: They emphasize the importance of continuously refining and optimizing their foundational models. This includes enhancing the model’s ability to process long contexts, accelerating inference speeds, and implementing other improvements that directly impact model performance and accuracy.
@@
4. AI search startups frequently opt for either open-source or commercialized big models as their foundational platform. They then integrate with search engine APIs, meticulously fine-tuning and optimizing these models to cater to the unique requirements of local markets and individual user preferences. Technology integration: AI search startups integrate advanced technologies such as semantic analysis and knowledge graphs into their foundational models. These integrations enable a deeper understanding of user queries, allowing for more relevant and accurate search results. Application scenarios: You can focus on specific domain or scenarios, such as e-commerce search, academic search, legal search, etc., or you can search in general domain. Optimization strategy: Continuously improve product performance and user experience through technological innovation and iterative optimization, while exploring cooperation opportunities with upstream and downstream enterprises.
Question: What is the relationship between graph computing, graph neural networks, and knowledge graphs? In the field of graph computing, do you think that more emphasis should be placed on research on graph database storage rather than graph neural networks? In addition, I would like to know what is the purpose of representing nodes as vectors in graph neural networks?
Graph computing, graph neural networks (GNNs), and knowledge graphs are closely related in AI and big data, but have distinct research focuses and applications.
Graph computing focuses on processing and analyzing graph-structured data, using algorithms like path search, centrality measures, and community detection. It encompasses not just graph database storage but also efficient data handling, feature extraction, and various computational tasks.
GNNs are deep learning models tailored for graph data. They capture dependencies through message passing and excel in tasks like classification, clustering, and link prediction. Their ability to handle non-Euclidean data makes them useful in social networks, recommendation systems, and bioinformatics. GNN research focuses on designing effective neural architectures and leveraging complex patterns for prediction and reasoning.
Knowledge graphs represent knowledge as graphs of entities and relationships, supporting applications like QA, recommendations, and search. They provide rich graph data for GNNs, which enhance knowledge graph representation and applications through tasks like node classification, link prediction, and relation reasoning.
In summary, knowledge graphs create graph data, graph computing explores it, and GNNs leverage it to enhance representation and application capabilities.
Graph database storage is a crucial aspect of graph computing, focusing on efficient and reliable storage of graph-structured data for subsequent processing and analysis. However, graph computing research extends beyond storage to encompass data handling, feature extraction, and various computational tasks on graphs.
Node vector representations form the basis for GNNs across tasks, capturing graph data’s intrinsic structure and patterns to aid prediction and reasoning. Without them, machine learning model training is unfeasible. Dense representations power deep learning, and enable efficient parallel computation, speeding up inference.
Question: With the emergence of big model technologies, is the relevance of knowledge graphs diminishing or being abandoned?
The rise of big model technology has profoundly impacted the construction and application of knowledge graphs. Despite their distinct architectures and application scopes, they complement each other, forming a symbiotic relationship. This synergy creates a new mode of dual knowledge engines (big models and knowledge graphs) working together to push the boundaries of AI development.
Big models, with their powerful representation learning, cross-domain generalization, and complex task-handling capabilities, breathe new life into AI. Trained on vast data, they capture richer, nuanced knowledge representations, addressing knowledge graphs’ limitations in coverage, update speed, and reasoning flexibility.
Meanwhile, knowledge graphs, as the cornerstone of structured knowledge storage and reasoning, offer precise, explainable, and easily queryable knowledge representations crucial for AI applications requiring high accuracy and interpretability. Their entities, relationships, and attributes provide a solid semantic foundation for intelligent services, enabling complex queries, reasoning, and decision support, mitigating big models’ hallucinations and logical errors.
Nowadays, knowledge graph projects are no longer isolated but deeply integrated with big model technology, forming a new paradigm of dual knowledge engines working in harmony. This collaboration enhances knowledge integration, utilization, and overall intelligence capabilities, such as understanding, reasoning, and creativity. big models can learn from knowledge graphs to improve domain-specific performance, while knowledge graphs can update and expand with new insights from big models.
Once a research hotspot, knowledge graphs have shown immense potential in information retrieval, question answering, and recommendation systems. With big model technology’s growth, knowledge graphs’ application scenarios have expanded further, jointly driving AI innovation.
In summary, the fusion of big model technology and knowledge graphs is an inevitable trend and a key to AI’s leapfrog development. As technologies mature and applications deepen, the dual knowledge engine model will play a pivotal role in various fields, fostering AI’s comprehensive development and widespread adoption.
@@ 5 Brain Science
On the other hand, the fruitful results of brain science [10] research are profoundly affecting the development trajectory of the field of artificial intelligence. By integrating the profound insights of brain science into the research and development of AI technology, it not only gives artificial intelligence systems capabilities that are closer to human thinking, but also greatly promotes the expansion of the boundaries and performance leaps of AI technology. This interdisciplinary integration not only provides a solid theoretical foundation and source of inspiration for the technological innovation of the AI industry, but also paves the way for the infinite possibilities of future intelligent technology.
@@
2. Memory mechanism: The human brain has a complex and sophisticated memory system, including short-term memory and long-term memory, as well as an efficient memory storage and retrieval mechanism. This biological characteristic provides valuable inspiration for the model architecture in memory processing, by drawing on the brain’s memory mechanism.
3. Multi-brain region collaborative information processing paradigm: Human cognitive functions do not exist in isolation, but rely on close collaboration and information exchange between multiple brain regions to form brain circuits (brainconnectome, brain connectivity) for cognitive functions. This multi-brain region collaborative mechanism helps the model’s design ideas when building complex information processing systems. By simulating the functional division of labor and synergy of different brain regions, more complex cognitive tasks can be achieved.
4. Dynamic system perspective of brain-inspired mechanism: As a highly dynamic system, the internal mechanism of the brain is far beyond the scope of simple electrical signal transmission. Complex phenomena such as the formation and forgetting of memories, the fluctuation and regulation of emotions, etc. often involve complex reactions and regulation of chemical substances. This perspective prompts us to not only focus on the optimization of the computational level when designing the Transformer model, but also explore how to introduce more diversified mechanisms (such as dynamic weight adjustment, emotional computing, etc.) to build a brain-inspired model that is closer to human intelligence.
5. Energy consumption: The energy consumption of the brain is much lower than that of the Transformer big model, mainly because: Efficient biological components: The energy efficiency of neurons and synapses far exceeds that of electronic components. Parallel and distributed processing: Information processing is highly parallel and distributed, which improves efficiency and reduces energy consumption. Sparse connections: The connections between neurons are sparse, which reduces unnecessary information processing. Adaptive plasticity: It can optimize neural networks based on learning and experience to reduce energy consumption. Evolutionary optimization: Long-term evolution has enabled the brain to develop efficient energy consumption mechanisms. Efficient energy utilization: Relying on glucose for energy, without energy reserves, emphasizing the importance of efficient energy consumption.
@@
The memory of agents can indeed be inspired by brain science. The brain’s memory mechanisms include short-term memory, long-term memory, working memory, etc. These mechanisms provide important references for the memory design of agents. For example, inspired by the brain’s working memory mechanism, the DNC (Differential Neural Computer) in artificial neural networks divides sequence control and memory storage into two modules, improving the ability to handle complex tasks. In addition, the continual learning mechanism in the brain also provides inspiration for agents to learn another new skill based on mastering one skill.